OpenStudy (anonymous):

what is a solution of the equation y=2x+5 ? (1point ) A. (5,20) B.(6,12) C.(3.11) D.(4,14)

OpenStudy (anonymous):

If left side = right side given the coordinates (x,y) from your solutions that will be your answer.

OpenStudy (anonymous):

is it C ? @iambatman

OpenStudy (aaliyahfairgood):

c

OpenStudy (aaliyahfairgood):

ill show you why

OpenStudy (aaliyahfairgood):

c.11=2(3)+5 11=6+5 11=11

OpenStudy (aaliyahfairgood):

always make sure the function is equal to each other